# Know your exceptions

### 01 Mar 2024 v1

@@ 1 Introduction
Imagine to be a modeller who is developing an ontology to represent knowledge about the zoological domain. Part of your knowledge are the propositions “birds fly” and “carnivora eat meat”. But, you know also that Tweety is a bird who does not fly (in fact she is a penguin) and that Po is a carnivorous who does not eat meat (because he is a panda). The question now is: how can we model this knowledge in our ontology? In fact, it seems reasonable wanting to not reject the information about birds and carnivora, since they are useful to provide knowledge regarding many individuals belonging to those classes. Therefore, we would like to model both the general properties of classes and the particular knowledge regarding the atypical individuals.
@@
The study of defeasible reasoning in the context of Artificial Intelligence (AI) led to the development of many logic-based accounts with the goal of modelling this kind of reasoning, called non-monotonic logics. “Non-monotonic” means exactly that adding new knowledge may lead to retract some conclusions, or, in other words, that some conclusions may not hold in all possible circumstances. However, these accounts rely on different intuitions and assumptions and so they may behave differently when it comes to their application to a specific domain. More specifically, they ‘behave differently’ for example in the way they spell out notions of ‘possible circumstance’ and thus what specific conclusions might or might not be supported by the formal system. This means that, without a making those intuitions and assumptions explicit, it is difficult to choose the system that suits best from an ontological point of view the specific domain the modeller is interested in capturing.
@@
In the classical, pre-Carnapian view of logic, it was often assumed that formal logic does not make any ontological commitments [11, 12]. Even speaking of ontological commitments in the case of admitting different logics may be considered implausible when we assume that each logic is agnostic with respect to the specific domain it is used to model and instead the choice of logic is merely a question of expressivity and formalisation style [13, 14]. Against this, we here hold that through the choices made in developing a non-monotonic formalism, a theory about defeasibility and exceptionality is conveyed and endorsed. Therefore, in order to make the ontological commitments explicit, and to understand the consequences of such choices, we need firstly a framework which allows us to extract and formulate such a foundational theory.
The central notions in our proposed framework are defeasibility and exceptionality. In particular, our analysis of the four systems for defeasibility will be guided by through an articulation of the ways in which these two properties are represented and interact. Thus, we start by developing a definition of “exception” in Section 2. In Section 3, we discuss the relation between exceptions and defeasibility through the introduction of the key distinction between universal generalisations and generic generalisations. In Section 4, we introduce briefly the four defeasible systems in order to provide the basis for the comparison in Section 5. We conclude with Section 6 by summarising the results and discussing some future directions for the research.
@@ 2 What Is an Exception?
In this section we discuss what does it mean to be an exception and propose a definition of this notion. We start with a distinction of different types of knowledge that explain why we can have exceptional entities, or in other terms different sources of defeasibility for knowledge.
@@ 2.1 Sources of Defeasibility
Before introducing the actual distinction it is important to note that here the concept of “knowledge” is used in a non-strict and technical way. In fact from a philosophical point of view, talking of ‘defeasible knowledge’ is contradictory since knowledge cannot be false in any sense. Therefore, we should speak of justified belief rather than full-fledged knowledge: however, for the sake of simplicity we prefer to use the term ‘knowledge’ even if we ask to keep in mind these important observation.
The first reason for having exceptions to something we know is that the knowledge we have is incomplete. In this case, we do not have enough knowledge to explain away all the exceptions. In this case we have a problem of completeness, we lack elements that could make the general picture coherent.
Another characteristic of our knowledge that makes it defeasible is uncertainty. In situations where knowledge is uncertain, we accept that there may be circumstances in which what we know does not really apply. In this case, the problem is justification, that is the fact that we have reasons justifying our knowledge which support it but do not affirm completely.
Different from uncertain knowledge is vague knowledge. The difference is that in the case of uncertain knowledge we are able to clearly distinguish those circumstances in which our knowledge correctly apply and those in which it does not. In the presence of vagueness, instead, we are not able to fix precisely the circumstances to which what we know apply. Here the issue is that of meaning: the borders of a concept are not clearly fixed.
The last case we consider of knowledge admitting exceptions is that of simplified knowledge. Differently from incomplete knowledge, we may know that precisely speaking there are not exceptions, but for the sake of simplicity we accept them. In other words, we may be able to make our knowledge precise enough that it encompasses also those cases which at first glance seem exceptions, however it would make the knowledge too complex and so we prefer to have a less precise but simpler version of it. In this case the problem is the precision, we accept less precision for the sake of simplicity.
These four types of knowledge are not mutually exclusive, for example we can have knowledge which is both uncertain and incomplete. However, this distinction can help in individuating the contexts in which we need to pay attention because we can have defeasible knowledge and so exceptions.
@@ 2.2 Common-Sense Definitions of “Exception”
In order to begin a delimitation of the concept of exception, we collected four common-sense definitions that we can find in vocabularies.
Cambridge dictionary: [^2] https://dictionary.cambridge.org/dictionary/english/exception
@@
Merriam-Webster dictionary: [^4] https://www.merriam-webster.com/dictionary/exception
“the act of excepting”, except (verb) [^6] https://www.merriam-webster.com/dictionary/excepting “to take or leave out from a number or a whole”;
Oxford languages: [^7] https://www.oed.com/dictionary/exception_n
@@
Collins dictionary: [^9] https://www.collinsdictionary.com/dictionary/english/exception
@@
All the above definitions rely on two main concepts for defining what an exception is: some kind of generalisation, expressed through terms like “rule”, “group”, “list”, “number”, “whole”, “general statements”; and an instance or individual which does not fit to or is excluded from that generalisation. Even if we endorse setting our definition in these terms, we think that we need to refine these definitions.
@@ 2.3 Exceptions, Counter-Examples and Errors
To operate this refinement we propose to analyse how the relation between the generalisation and the individual can take place. First of all, what we need to stress is the fact that the exclusion of the individual happens in spite of the fact that we expect it to be included. For example, if we are speaking of flying birds, a dog cannot be considered an exception, since we do not expect it to satisfy that generalisation.
However, there is still the possibility that we are not speaking of exceptions even in that case. In fact, assume that we have a generalisation and an instance to which that generalisation should apply, but it actually does not, that is an instance which is excluded from that generalisation. This discrepancy can happen in three cases:
  • Counter-example: the generalisation is falsified by the instance, e.g. The school can be reached only by bus or train, but Bob arrives by car.
  • Error: the instance is faulty, e.g. Students have an unique ID card number, but Alice is registered with a wrong, non-unique number.
  • Exception: both the generalisation and the instance are correct, e.g. Students have no salary, but Eve is a PhD student with a bursary.
As we can see, according to the truth-values we give to the propositions describing the generalisation and the instance we have three different situations. The peculiarity of the case of exceptions is that even if the two propositions contradict each other, we want to consider both of them true. The reason for this can be recognised in the kind of knowledge we are using, that is incomplete, uncertain, vague or simplified knowledge. However, this means that we cannot rely on classical deductive reasoning, since this would lead to a contradiction.
So, now we collected the pieces that we can use to elaborate a definition of exceptions. Firstly, we need to be in a particular context where the knowledge we have is incomplete, uncertain, vague or simplified. Secondly, we need two elements for individuating exceptions, a generalisation and an individual that for some reason is excluded from that generalisation, even if it should be included. Finally, we want to consider both the knowledge of the generalisation and of that instance as correct. Therefore, we give our explication of the notion of ‘exception’ as follows:
@@
This definition is our starting point for the discussion in the next sections, which will lead to more refined version of it.
@@ 3 Defeasibility and Exceptions
To start refining the definition of exceptions we can start from one of the main elements that compose it, that is the notion of generalisation. We begin the analysis by looking at some relevant fields where generalisations that admit exceptions have been studied.
@@ 3.1 Defeasible Generalisations, Generics and Ceteris Paribus Laws
We can start from the distinction made in [2] between universals and generics, even if slightly re-elaborated: if in [2] those terms are referred to properties, we will instead use them for propositions in order to be coherent with the working-definition 2.3 developed above. In particular, universal propositions correspond to generalisations like Dogs are animals, that is where the predicated property applies to all the instances of the first class, whereas generic propositions correspond to those like our examples above, where there could be instances that do not instantiate the predicated property, like Po the panda bear for Carnivora eat meat.
@@ 3.1.1 Generics
An interesting insight on this distinction comes from the literature on generics [15, 16]. Generics are defined as sentences reporting a regularity regarding particular facts that can be generalised even if they admit exceptions. In the penguin example, for instance, the sentence birds fly can be considered a generic since it states something about the individuals of the class bird, but still allowing exceptions like the members of the sub-classes penguins, ostriches and emus. The other peculiarity beyond tolerating exceptions is that in this kind of generalisations we do not have information about how many individuals satisfy the generalisation.
Since we are interested in generalisations that admit exceptions in general, included those indicating quantities like most books are paperback or no car is able to fly, we do not commit to considering generic generalisations. For this reason we will refer to those generalisations more generally as defeasible generalisations, highlighting their capacity to admit exceptions and so fail to apply to all the cases. However, the research of linguists, philosophers and cognitive scientists in generics, suggest the distinction between universal generalisations and defeasible ones may not be only a methodological distinction but that they are really distinct phenomena, at least in the sub-kind of generics.
@@ 3.1.2 Ceteris paribus laws
Another related topic is that of ceteris paribus laws (henceforth cp-laws), which attains the general fields of philosophy of science and metaphysics. Considering special sciences like biology, psychology or economics, it has been noted that most of the laws they state are not universal. That is, in special sciences “there are (actual and merely possible) situations in which the above generalizations do not hold, although all the conditions obtain that are explicitly stated in the antecedents of these generalizations” [17]. These laws have been named ceteris paribus, which is a Latin terminology that can be translated as “all other things been equal”: as we can read from the above quotation, such laws are strictly related to the notions we are considering in our analysis on defeasibility in terms of generalisations and exceptions. The idea behind cp-laws is that they are valid assuming that other factors that may influence their application to a specific case remain constant or are completely absent. Determining the precise meaning of this is the goal of studies on cp-laws: for the time being we are more interested in some insights we can gain for our purposes from such studies.
@@
First of all, we note that the generalisations as we have considered them until now can be seen as weak laws we apply everyday in common-sense reasoning: in this sense, cp-laws as laws of the special sciences are in the mid-way between common-sense generalisations and the generalisations made for example in fundamental physics, which are considered paradigmatic universal laws. For this reason, the insights coming from the study of cp-laws with respect of their peculiar capability of admitting exceptions can a fortiori be applied to defeasible generalisations which are weaker generalisations not refined by the process of scientific inquiry.
Definite and indefinite ceteris paribus laws.
@@
The first interesting observation regards exclusive cp-laws, that is those for which disturbing factors have to be absent, which can be distinguished between definite and indefinite cp-laws [17, section 3.2]. The former are those for which the disturbing factors excluded from the antecedent are specified, that is, it is possible to complete the cp-law with those factors. Intuitively, indefinite cp-laws are those for which it is not possible to do such a completion if not in the form of adding to the antecedent the clause “excluding all the disturbing factors” without specifying what these disturbing factors are.
@@
We return on this point in the next section because the possibility to complete the cp-law suggest that it may be possible to do the same for defeasible generalisations and this allows to transform a defeasible generalisation in an equivalent universal generalisation through the addition of this complement.
The dilemma.
@@
The last observation comes exactly from the possibility of the transformation suggested above. In fact, exclusive cp-laws are subject to a dilemma [17, section 4]: if we transform a cp-law in a universal law, it will tend to be false, since it is implausible that all the interfering factors have been considered and even only one exception would turn into a counter-example to a universal law. Note that a similar observation is done in [2] with respect to the strategy of listing all the exception to a generalisation.
@@
On the other side, if we add a completer like that for indefinite exclusive cp-laws, that is similar to “assuming that nothing interferes”, the risk is to end up with a trivially true statement given its unwelcome similarity to a statement like “All As are Bs or not all As are Bs’. This is particularly interesting in the case of a modeller, since even if we avoid to transform a defeasible generalisation into a tautology, it would remain very little informative about the actual conditions under which the generalisation should apply and under which it should admit exceptions instead.
@@
In the next section we will discuss the relationship between universal and defeasible generalisations in light of the features discussed above and especially of the dilemma emerging for cp-laws.
@@ 3.2 Universal and defeasible generalisations
Since what characterises the generalisation is the quantifier, we can represent defeasible generalisations by substituting the quantifier with something different. Therefore, we denote defeasible generalisations with the symbol $\widetilde{\forall}$ and refer to them with the term $\widetilde{\forall}$-generalisations. [^11] Note that a similar move is done to represent syntactically generics: the quantifier Gen is introduced for generic sentences [15]. Because we do not want to commit ourselves with generics specifically, we will not use Gen, but instead use the symbol $\widetilde{\forall}$. Consequently, the aim of non-monotonic logics can be seen as introducing a formal way to model $\widetilde{\forall}$-generalisations. As we have said above, there are many ways this has been done, and we will analyse them below. A first naive intuition that can be adopted is listing all the exceptional cases and excluding them [2]. However, since most $\widetilde{\forall}$-generalisations are like indefinite cp-laws for which it is not possible to be sure that all the exceptions have been considered, this is not a good strategy and, as we mentioned above, we are exposed to the first horn of the dilemma 3.1.2.
So, as [2] points out, we end up saying something like Carnivora eat meat except those that do not, and the real goal becomes characterising somehow “those that do not”. In other words, the goal becomes to characterise the class of exceptions without falling on the second horn of the dilemma, since the impression is that now the defeasible generalisation is trivial. Nevertheless, this problem gives us a first insight on a relation between $\forall$-generalisations and $\widetilde{\forall}$-generalisations.
In fact, the idea here is that a defeasible generalisation can be seen as a universal generalisation where some instances are removed from the generalisation, namely the exceptions to that generalisation considered as universal. Consequently, we can exploit this intuition to define $\widetilde{\forall}$-generalisations:
Definition 2
@@
For instance, consider the case of the statement: Prime numbers are odd numbers. There is a clear exceptions which is the prime number 2. We can apply the equation above and write “~$\widetilde{\forall}x(x\ \text{is a prime number}\rightarrow x\ \text{is an oddnumber})\Longleftrightarrow\forall x(x\ \text{is a prime number}\rightarrow x\ \text{is an odd number})-\{2\}$”. This was a simple case where we know for sure that there is only one exception, however, we can see that in more complex cases the key point of $\widetilde{\forall}$-generalisations represented as universal ones is how to characterise the class of exceptions. In fact, according to the definitions of these two kinds of generalisations we can see that the complement of the subset identified by the quantifier $\widetilde{\forall}$ is the set of the exceptions. In this sense the second horn of dilemma 3.1.2 should be made precise by saying that a formulation like the second term in the equivalence is trivial only if we are not able to individuate the exceptions to that generalisation.
@@
Related to this, another observation is relevant: the class of exceptions is relative to the generalisation we are considering and we cannot possibly have exceptions without a generalisation. Consequently, in principle, it cannot be defined a priori, that is before the statement of the generalisation. For instance, in the case of Tweety, you cannot exclude from the class of birds the exceptional ones, without firstly saying that we are considering the defeasible generalisation Birds fly.

### 05 Mar 2024 v2

@@ 1 Introduction
Imagine being a modeller who is developing an ontology to represent knowledge about the zoological domain. Part of your knowledge are the propositions “birds fly” and “carnivora eat meat”. But, you know also that Tweety is a bird who does not fly (in fact she is a penguin) and that Po is a carnivorous who does not eat meat (because he is a panda). The question now is: how can we model this knowledge in our ontology? In fact, it seems reasonable wanting to not reject the information about birds and carnivora, since they are useful to provide knowledge regarding many individuals belonging to those classes. Therefore, we would like to model both the general properties of classes and the particular knowledge regarding the atypical individuals.
@@
The study of defeasible reasoning in the context of Artificial Intelligence (AI) led to the development of many logic-based accounts with the goal of modelling this kind of reasoning, called non-monotonic logics. “Non-monotonic” means exactly that adding new knowledge may lead to retract some conclusions, or, in other words, that some conclusions may not hold in all possible circumstances. However, these accounts rely on different intuitions and assumptions and so they may behave differently when it comes to their application to a specific domain. More specifically, they ‘behave differently’ for example in the way they spell out notions of ‘possible circumstance’ and thus what specific conclusions might or might not be supported by the formal system. This means that, without making those intuitions and assumptions explicit, it is difficult to choose the system that suits best from an ontological point of view the specific domain the modeller is interested in capturing.
@@
In the classical, pre-Carnapian view of logic, it was often assumed that formal logic does not make any ontological commitments [11, 12]. And even when, following Carnap, we admit different logics, speaking of ontological commitments may still be considered implausible if we assume that each logic is agnostic with respect to the specific domain it is used to model and thus the choice of logic is merely considered a question of expressivity and formalisation style [13, 14]. Against this, we here hold that through the choices made in developing a non-monotonic formalism, a theory about defeasibility and exceptionality is conveyed and endorsed. Therefore, in order to make the ontological commitments explicit, and to understand the consequences of such choices, we need firstly a framework which allows us to extract and formulate such a foundational theory.
The central notions in our proposed framework are defeasibility and exceptionality. In particular, our analysis of the four systems for defeasibility will be guided by an articulation of the ways in which these two properties are represented and interact. Thus, we start by developing a definition of “exception” in Section 2. In Section 3, we discuss the relation between exceptions and defeasibility through the introduction of the key distinction between universal generalisations and defeasible generalisations. In Section 4, we introduce briefly the four defeasible systems in order to provide the basis for the comparison in Section 5. We conclude with Section 6 by summarising the results and discussing some future directions for the research.
@@ 2 What Is an Exception?
In this section, we discuss what it means to be an exception and propose a definition of this notion. We start with a description of sources of defeasibility for knowledge [^2] Note that here the concept of “knowledge” is used in a non-strict and non-technical way. In fact, from a philosophical point of view, talking of “defeasible knowledge” is contradictory since knowledge cannot be false by the very definition of the term, e.g. in the tradition explicating knowledge as variants of “justified true belief”. Therefore, we should speak of justified belief rather than full-fledged knowledge: however, for the sake of simplicity, we prefer to use the term “knowledge” even if we ask to keep in mind this important observation., that is, a distinction between different types of knowledge that are used to explain why an entity may be considered exceptional.
@@ 2.1 Sources of Defeasibility
The first reason for having exceptions to something we know is that the knowledge we have is incomplete: we do not have enough knowledge to explain away all the exceptions. In this case, we have a problem of completeness, we lack elements that could make the general picture coherent.
Another feature of our knowledge that makes it defeasible is uncertainty. In situations where knowledge is uncertain, we accept that there may be circumstances in which what we know does not really apply. In this case, the problem is justification, that is, the fact that we have reasons justifying our knowledge which support it but do not affirm it completely. Different from uncertain knowledge is vague knowledge. The difference is that in the case of uncertain knowledge we are able to clearly distinguish those circumstances in which our knowledge correctly applies and those in which it does not. In the presence of vagueness, instead, we are not able to fix precisely the circumstances in which what we know applies. Here the issue is that of meaning: the borders of a concept are not clearly fixed. The last case we consider of knowledge admitting exceptions is that of simplified knowledge. Differently from incomplete knowledge, we may know that precisely speaking there are no exceptions, but for the sake of simplicity we accept some as such. In other words, we may be able to make our knowledge precise enough that it encompasses also those cases which at first glance seem exceptions, however it would make the knowledge too complex and so we prefer to have a less precise but simpler version of it. In this case the problem is the precision, we accept less precision for the sake of simplicity.
These four types of knowledge are not mutually exclusive, for example we can have knowledge which is both uncertain and incomplete. However, these distinctions will help us in identifying those contexts in which we need to pay particular attention to the possibility of defeasible knowledge, and thus to the presence of exceptions.
@@ 2.2 Common-Sense Definitions of “Exception”
In order to begin a delimitation of the concept of exception, we collected four common-sense definitions for the term “Exception” that we can find in dictionaries.
Cambridge dictionary:33footnotemark: 3
@@
Oxford languages:55footnotemark: 5
@@
Collins dictionary:77footnotemark: 7
@@
All the above definitions rely on two main concepts for defining what an exception is: some kind of generalisation, expressed through terms like “rule”, “group”, “list”, “general statements”; and an instance or individual which does not fit to or is excluded from that generalisation. Even if we endorse the approach of setting up our definition in these terms, we clearly need to ontologically refine the underlying ideas.
@@ 2.3 Exceptions, Counter-Examples and Errors
To carry out our refinement of this core idea, we propose to analyse how the relation between the generalisation and the individual can take place. First of all, what we need to stress is the fact that the exclusion of the individual happens in spite of the fact that we expect it to be included. For example, if we are speaking of flying birds, a dog cannot be considered an exception, since we do not expect it to satisfy that generalisation.
However, there is still the possibility that we are not speaking of exceptions even in that case. In fact, assume that we have a generalisation and an instance to which that generalisation should apply, but it actually does not, that is, it is an instance which is excluded from that generalisation. This discrepancy can happen in three cases:
Counter-example:
the generalisation is falsified by the instance, e.g.:
The school can be reached only by bus or train, but Bob arrives by car.
Error:
the instance is faulty, e.g.:
Students have a unique ID card number, but Alice is registered with a wrong, non-unique number.
Exception:
both the generalisation and the instance are correct, e.g.:
Students have no salary, but Eve is a PhD student with a bursary.
As we can see, according to the truth-values we give to the propositions describing the generalisation and the instance we have three different situations. The peculiarity in the case of exceptions is that even if the two propositions contradict each other, we want to consider both of them true. The reason for this can be recognised in the kind of knowledge we are using, namely whether it is incomplete, uncertain, vague or simplified knowledge. However, this means that we cannot rely on classical deductive reasoning, since this would lead to a contradiction.
We now collected the basic pieces that we can use to elaborate a definition of exceptions. Firstly, we need to be in a particular context where the knowledge we have is incomplete, uncertain, vague or simplified. Secondly, we need two elements for individuating exceptions, a generalisation and an individual that for some reason is excluded from that generalisation, even if it should be included. Finally, we want to consider both the knowledge about the generalisation and the particular individual as correct. Therefore, we give our explication of the notion of ‘exception’ as follows:
@@
This explication is our starting point for the discussion in the next sections, which will lead to a more refined version of this position.
@@ 3 Defeasibility and Exceptions
To begin refining the definition of exceptions we can start from one of the main elements that compose it, that is the notion of generalisation. We begin the analysis by looking at some relevant fields where generalisations that admit exceptions have been studied.
@@ 3.1 Defeasible Generalisations, Generics and Ceteris Paribus Laws
We can start from the distinction made in [2] between universals and generics, even if slightly re-elaborated: if in [2] those terms are used for properties, we will instead use them for propositions in order to be coherent with the working definition developed above in Section 2.3. In particular, universal propositions correspond to generalisations such as Dogs are animals, that is, where the predicated property applies to all the instances of the first class, whereas generic propositions correspond to those like our examples above, where there could be instances that do not instantiate the predicated property, such as Po the panda bear for Carnivora eat meat.
@@ 3.1.1 Generics
An interesting insight regarding this distinction comes from the literature on generics [15, 16]. Generics are defined as sentences reporting a regularity regarding particular facts that can be generalised even if they admit exceptions. In the penguin example, for instance, the sentence Birds fly can be considered a generic since it states something about the individuals of the class bird, but still allows exceptions like the members of the sub-classes penguins, ostriches and emus. The other peculiarity beyond tolerating exceptions is that in this kind of generalisations we do not have information about how many individuals satisfy the generalisation.
Since we are interested in generalisations that admit exceptions in general, including those indicating quantities like most books are paperback or no car is able to fly, we do not restrict ourselves to considering just generic generalisations. For this reason, we will refer to those generalisations more broadly as defeasible generalisations, highlighting their capacity to admit exceptions and thus fail to apply to all the cases. However, research of linguists, philosophers and cognitive scientists in generics suggests that the distinction between universal generalisations and defeasible ones may not be only a methodological distinction but that they are really distinct phenomena, at least in the sub-kind of generics [17].
@@ 3.1.2 Ceteris paribus laws
Another related topic is that of ceteris paribus laws (henceforth cp-laws), which attains the general fields of philosophy of science and metaphysics. Considering scientific fields such as biology, psychology or economics, it has been noted that most of the laws they state are not universal. That is, in particular sciences “there are (actual and merely possible) situations in which the above generalisations do not hold, although all the conditions obtain that are explicitly stated in the antecedents of these generalizations” [18]. These laws have been termed ceteris paribus, which is a Latin expression that can be translated as “all other things being equal”: as we can read from the above quotation, such laws are strictly related to the notions we are considering in our analysis on defeasibility in terms of generalisations and exceptions. The idea behind cp-laws is that they are valid assuming that other factors that may influence their application to a specific case remain constant or are completely absent. Determining the precise meaning of this is the goal of studies on cp-laws: for the time being, we are more interested in some insights we can gain for our purposes from such studies.
@@
First of all, we note that the generalisations as we have considered them until now can be seen as weak laws we apply everyday in common-sense reasoning: in this sense, cp-laws as laws of the particular sciences are in the mid-way between common-sense generalisations and the generalisations made, for instance, in fundamental physics, which are considered paradigmatic universal laws. For this reason, the insights coming from the study of cp-laws with respect of their peculiar capability of admitting exceptions can a fortiori be applied to defeasible generalisations which are weaker generalisations not refined by the process of scientific inquiry.
@@
A first interesting observation concerns exclusive cp-laws, that is, those for which interfering factors have to be absent, which can be distinguished between definite and indefinite cp-laws [18, Section 3.2]. The former are those for which the interfering factors excluded from the antecedent are specified, that is, it is possible to complete the cp-law with those factors. Intuitively, indefinite cp-laws are those for which it is not possible to do such a completion if not in the form of adding to the antecedent the clause “excluding all the interfering factors” without specifying what these interfering factors are.
@@
We return to this point in the next section because the possibility of completing the cp-law suggests that it may be possible to do the same for defeasible generalisations and this allows to transform a defeasible generalisation into an equivalent universal generalisation through the addition of this complement.
@@
The last observation comes exactly from the possibility of the transformation suggested above. In fact, exclusive cp-laws are subject to a dilemma [18, Section 4]: if we transform a cp-law into a universal law, it will tend to be false, since it is implausible that all the interfering factors have been considered and even only one exception would turn into a counter-example to a universal law. Note that a similar observation is raised in [2] with respect to the strategy of listing all the exceptions to a generalisation.
@@
On the other side, if we add a completer like that for indefinite exclusive cp-laws that is similar to “assuming that nothing interferes”, the risk is to end up with a trivially true statement given its unwelcome similarity to a statement like “All As are Bs or not all As are Bs”. This is particularly interesting from the perspective of a modeller, since even if we avoid transforming a defeasible generalisation into a tautology, it would remain very moderately informative about the actual conditions under which the generalisation should apply and under which it should admit exceptions instead.
@@
In the next section, we will discuss the relationship between universal and defeasible generalisations in light of the features discussed above and especially of the dilemma emerging for cp-laws.
@@ 3.2 Universal and defeasible generalisations
Since what characterises the generalisation is the quantifier, we can represent defeasible generalisations by substituting the quantifier with something different. Therefore, we denote defeasible generalisations with the symbol $\widetilde{\forall}$ and refer to them with the term $\widetilde{\forall}$-generalisations. [^9] Note that a similar move is done to represent syntactically generics: the quantifier Gen is introduced for generic sentences [15]. Because we do not want to commit ourselves with generics specifically, we will not use Gen, but instead use the symbol $\widetilde{\forall}$. Consequently, the aim of non-monotonic logics can be seen as introducing a formal way to model $\widetilde{\forall}$-generalisations. As we have said above, there are many ways this has been done, and we will analyse them below. A first naive intuition that can be adopted is listing all the exceptional cases and excluding them [2]. However, since most $\widetilde{\forall}$-generalisations are like indefinite cp-laws for which it is not possible to be sure that all the exceptions have been considered, this is not a good strategy and, as we mentioned above, we are exposed to the first horn of the dilemma LABEL:dilemma.
So, as [2] points out, we end up saying something like Carnivora eat meat except those that do not, and the real goal becomes characterising somehow “those that do not”. In other words, the goal becomes to characterise the class of exceptions without falling on the second horn of the dilemma, since the impression is that now the defeasible generalisation is trivial. Nevertheless, this problem gives us a first insight into a relation between $\forall$-generalisations and $\widetilde{\forall}$-generalisations.
In fact, the idea here is that a defeasible generalisation can be seen as a universal generalisation where some instances are removed from the generalisation, namely the exceptions to that generalisation considered as universal. Consequently, we can exploit this intuition to precisify the meaning of $\widetilde{\forall}$-generalisations semi-formally:
Concept Explication. 2
@@
The meaning of “strict subset” is defined by each particular formal system.
@@
For instance, consider the case of the statement: Prime numbers are odd numbers. There is a clear exception which is the prime number 2. We can apply the equivalence above and write “~$\widetilde{\forall}x(x\ \text{is a prime number}\rightarrow x\ \text{is an oddnumber})\Longleftrightarrow\forall x(x\ \text{is a prime number}\rightarrow x\ \text{is an odd number})-\{2\}$”. This was a simple case where we know for sure that there is only one exception, however, we can see that in more complex cases the key point of $\widetilde{\forall}$-generalisations represented as universal ones is how to characterise the class of exceptions. In fact, according to the definitions of these two kinds of generalisations we can see that the complement of the subset identified by the quantifier $\widetilde{\forall}$ is the set of the exceptions. In this sense, the second horn of dilemma LABEL:dilemma should be made precise by saying that a formulation like the second term in the equivalence is trivial only if we are not able to individuate the exceptions to that generalisation.
@@
Related to this, another observation is relevant: the class of exceptions is relative to the generalisation we are considering and we cannot possibly have exceptions without a generalisation. Consequently, in principle, it cannot be defined a priori, that is before the statement of the generalisation. For instance, in the case of Tweety, you cannot exclude from the class of birds the exceptional ones, without first saying that we are considering the defeasible generalisation Birds fly.